### v1

Abstract
Accurately and swiftly assessing damage from conflicts is crucial for humanitarian aid and regional stability. In conflict zones, damaged zones often share similar architectural styles, with damage typically covering small areas and exhibiting blurred boundaries. These characteristics lead to limited data, annotation difficulties, and significant recognition challenges, including high intra-class similarity and ambiguous semantic changes. To address these issues, we introduce a pre-trained DINOv3 model and propose a multi-scale cross-attention difference siamese network (MC-DiSNet). The powerful visual representation capability of the DINOv3 backbone enables robust and rich feature extraction from bi-temporal remote sensing images. The multi-scale cross-attention mechanism allows for precise localization of subtle semantic changes, while the difference siamese structure enhances inter-class feature discrimination, enabling fine-grained semantic change detection. Furthermore, a simple yet powerful lightweight decoder is designed to generate clear detection maps while maintaining high efficiency. We also release a new Gaza-change dataset containing high-resolution satellite image pairs from 2023-2024 with pixel-level semantic change annotations. It is worth emphasizing that our annotations only include semantic pixels of changed areas. Unlike conventional semantic change detection (SCD), our approach eliminates the need for large-scale semantic annotations of bi-temporal images, instead focusing directly on the changed regions. We term this new task change semantic detection (CSD). The CSD task represents a direct extension of binary change detection (BCD). Due to the limited spatial extent of semantic regions, it presents greater challenges than traditional SCD tasks. We evaluated our method under the CSD framework on both the Gaza-Change and SECOND datasets. Experimental results demonstrate that our proposed approach effectively addresses the CSD task, and its outstanding performance paves the way for practical applications in rapid damage assessment across conflict zones.
keywords:
[…]
1 Introduction
Accurately and timely assessing damage zones in conflict areas is a critical task with profound implications for humanitarian assistance, disaster relief, and post-conflict reconstruction QING2022102899; HOLAIL202548. Similar to building damage assessment in natural disasters han2025multi, remote sensing images, particularly high-resolution satellite data, have become an indispensable tool for large-scale monitoring of these changes. However, whereas previous building damage assessments caused by natural disasters focused more on binary changes, damage assessment in conflict areas may place greater emphasis on fine-grained types. Therefore, the core task in conflict zones is semantic change detection (SCD) of buildings.
In recent years, the remarkable success of deep learning in both computer vision od2022review; sultana2020evolution and natural language processing guo2025deepseek has profoundly reshaped the landscape of remote sensing semantic change detection. The evolution of SCD methodologies has largely mirrored broader trends in visual recognition. The field initially adopted convolutional neural network (CNN)-based architectures, drawing direct inspiration from semantic segmentation tasks peng2019end. However, these early, multi-stage pipelines often suffered from significant error accumulation. This limitation prompted a shift towards siamese network structures, which emerged as a more robust foundational paradigm for direct change representation DLCD2018FC. Subsequently, with the rise of vision transformers (ViTs) vaswani2017attention; han2022survey; liu2021swin, transformer-based architectures have begun to establish a new state-of-the-art for SCD. Nevertheless, a fundamental challenge persists: these data-hungry deep models require vast amounts of meticulously annotated data, which is notoriously difficult and expensive to obtain for remote sensing applications.
To address the aforementioned challenges, we propose a new paradigm termed change semantic detection (CSD). Distinct from conventional semantic change detection (SCD), our approach eliminates the need for annotating entire semantic regions, instead focusing solely on change masks. This framework represents a direct extension of binary change detection (BCD). While significantly reducing annotation difficulty and time requirements, this new paradigm consequently increases the challenge of limited target region proportions. This necessitates substantial improvements in the model’s capability to extract features from small target areas.
Figure 1: Evolution of change detection paradigms: (a) Binary Change Detection (BCD), (b) Semantic Change Detection (SCD), and (c) our proposed Change Semantic Detection (CSD).
Based on the above analysis, we identify four major challenges in framing conflict-induced damage assessment as a CSD task: (1) Inherent data scarcity: Limited by the geographical extent of conflict zones and the number of destroyed areas available for training. (2) Small target regions: CSD focuses exclusively on damaged areas, resulting in minimal semantic region coverage. (3) Subtle and ambiguous changes: Infrastructure damage in conflict zones varies significantly in severity and extent, particularly making minor damage difficult to detect. (4) High inter-class similarity: Different facility categories within the same region may share similar characteristics, making fine-grained damage assessment particularly challenging for semantic change detection.
To bridge this gap, we draw inspiration from the recent success of foundational vision models. We argue that leveraging their rich, pre-trained representations is key to overcoming data scarcity and recognizing subtle semantic changes. In this paper, we propose a novel DINOv3-driven siamese network for CSD. Specifically, we adopt the DINOv3 simeoni2025dinov3 model pre-trained on satellite data, with ConvNeXt liu2022convnet as its main backbone architecture, which helps reduce the distribution discrepancy between the pre-training data and the actual application data. Then, we propose a multi-scale attention mechanism to extract and enhance features at different levels, aiming to capture the subtle and ambiguous change features of infrastructure damage. Furthermore, we perform an absolute value differential operation on the obtained semantic-rich feature maps to increase inter-class feature differences. Finally, a carefully designed decoder network with attention enhancement is used to generate clear semantic change detection maps. We also release a building semantic change detection dataset of the Gaza area from 2023 to 2024. As shown in Figure 2, we present panoramic remote sensing images of the Gaza Strip captured by satellites. To the best of our knowledge, this is the first remote sensing semantic change detection study focused on conflict area assessment, laying a foundation for future research in related fields. In summary, our work makes the following key contributions:
  • We introduce a multi-scale cross-attention difference siamese network (MC-DiSNnet). Built upon a pre-trained DINOv3 backbone, our network extracts robust, generalized features. The cross-attention mechanism is strategically employed to fuse multi-scale temporal features, enabling it to pinpoint subtle, semantic-changing regions effectively.
  • We contribute a new dataset for the Gaza area, containing high-resolution bi-temporal satellite image pairs from 2023-2024 with meticulously annotated pixel-level semantic change labels. To our knowledge, this is the first change detection study specifically focused on conflict area assessment.
  • We propose the novel change semantic detection (CSD) paradigm that fundamentally shifts from exhaustive bi-temporal semantic annotation to focused labeling of changed semantic regions. This strategic simplification significantly reduces annotation complexity and human labor.
  • Extensive experiments show that our method achieves state-of-the-art performance not only on our proposed Gaza dataset but also on the public benchmarks, SECOND yang2021asymmetric, demonstrating its superior robustness and generalization capability.
Figure 2: The panoramic remote sensing satellite image of the Gaza Strip.
2 Related Work
In this section, topics related to remote sensing image change detection (RSCD) are introduced, including methods based on classical AI approaches and those based on vision foundation models.
2.1 Classical Deep Learning for Change Detection
Change detection involves multiple inputs. Traditional methods typically require multiple processing stages. To avoid error accumulation effects, peng2019end pioneered the use of segmentation networks for change detection tasks. By pairing the input images, this single-stage approach not only shortens the detection pipeline but also enhances detection performance. Subsequently, models based on classical convolutions and Transformer models focusing on global information have emerged as key research foci in change detection. DLCD2018FC employed a siamese network for change detection, as its architecture is inherently suited to this task. On the other hand, DLCD2021transformer introduced a Transformer module after the CNN features to handle global information, addressing the challenges posed by high-resolution remote sensing images. WNet tang2023wnet builds upon this CNN-Transformer hybrid architecture by incorporating deformable convolutions and designing a feature fusion module (CTFM) to integrate local, global, and cross-scale features from CNN and Transformer encoders.
To achieve better fusion within such hybrid models, feng2022icif proposed a parallel architecture integrating CNN and Transformer to simultaneously capture local and global features. Specifically, an intra-scale cross-interaction module was designed to interact with convolutional and Transformer features, followed by an inter-scale feature fusion module for integration. In contrast, he2024cbsasnet devised a channel-bias separation attention (CBSA) module, which enhances the extraction of detailed information by integrating features from multiple receptive fields. changeformer introduced ChangeFormer, a novel architecture combining Transformer with multi-layer perceptron (MLP). This model effectively captures both spatial and temporal features, making it well-suited for remote sensing change detection tasks. Focusing on urban scenarios, zhan2025difference augmented multi-scale feature perception capabilities to detect urban building features of varying scales in remote sensing imagery.
2.2 Vision Foundation Models
Visual Foundation Models (VFMs) VFM represent a category of AI models built upon deep neural networks and self-supervised learning. Trained on large-scale, multi-source datasets, they excel at capturing complex visual features and demonstrate superior performance across diverse visual tasks, including object detection, image classification, and semantic segmentation. Through pre-training on vast datasets, VFMs acquire general recognition capabilities that readily transfer to various downstream tasks. Beyond achieving state-of-the-art results in standard supervised learning, VFMs also facilitate advanced functionalities like zero-shot and interactive segmentationsamclip.
The advent of VFMs has significantly transformed deep learning-based visual processing. Masked Autoencoders (MAE) pioneered a vision masking strategy utilizing Transformers, achieving outstanding performance by heavily masking large portions of input images. CLIP clip leverages massive image-text pairs for training, effectively boosting image-text similarity through metric learning. Segment Anything Model (SAM) sam and its derivatives (e.g., FastSAM fastsam, MobileSAM mobilesamv2) mark a paradigm shift in image segmentation, demonstrating remarkable proficiency in detecting, segmenting, and generating objects of any category, showcasing exceptional versatility.
For remote sensing scenarios, visual foundation models (VFMs) still require effective fine-tuning to achieve satisfactory performance. Pioneering the application of VFMs to remote sensing change detection, SAMCD proposed the SAM-CD architecture. This approach employs a lightweight version of FastSAM as the visual encoder and incorporates a convolutional adaptor to enhance its adaptability to remote sensing scenes. Building upon FastSAM, gao2025combining utilized fine-tuning adapters to reduce data dependency. Conversely, zhang2024integrating introduced DoRA (weight-decomposed low-rank adapter) for fine-tuning and designed a boundary distance-based loss function to optimize segmentation boundaries, specifically addressing the challenges of small objects and ambiguous edges. Focusing on lightweight performance, mobilesambased1 and SCDSAM developed efficient yet effective models based on MobileSAM.
3 Method
3.1 Overall Architecture
The proposed multi-scale cross-attention difference sia- mese network (MC-DiSNet), based on the DINOv3 simeoni2025dinov3 backbone, employs a two-branch encoder-decoder architecture consisting of three key components: an enhanced DINOv3 backbone, a multi-scale cross-attention difference (MSCAD) module, and a lightweight yet powerful decoder. The enhanced backbone preserves three hierarchical stages and injects change awareness using bottleneck adapters, multi-stage prompts, and LoRA.
The MSCAD module processes dual-branch features from the encoder. It aligns the tri-stage features with a cross-scale adapter, employs a diff-processor to capture temporal discrepancies, and integrates direct and adaptive differences via a diff-aggregator to produce sharp multi-scale change cues.
Subsequently, the decoder refines these cues: its depth-wise residual context enhancer reconstructs thin structures, and an attention gate filters out background noise. The result is a noise-robust, edge-accentuated feature map ready for final decoding.
Figure 3: Overall architecture of the proposed MC-DiSNet.
3.2 Enhanced DINOv3 Backbone
To maximize the feature extraction capability of the DINOv3 backbone while balancing computational cost and accuracy for the proposed change semantic detection (CSD) task, we introduce an enhanced DINOv3 encoder. This encoder keeps the original ConvNeXt-Tiny weights frozen and enriches multi-scale features C3-C5 through three lightweight plug-ins. First, a bottleneck adapter is inserted after each MLP block, comprising a down-projection, GELU activation lee2023mathematical, and an up-projection branch with a reduction ratio. Its forward pass is formulated as:
$\mathbf{y}=\mathbf{x}+s\cdot\mathrm{Up}\bigl(\mathrm{GELU}(\mathrm{Down}(\mathbf{x}))\bigr),$ (1)
where the scalar gate $s$ is initialized to zero and learned during training. Subsequently, 20 learnable prompt tokens are concatenated with the C2-C4 feature maps and updated via cross-stage projection. This steers the network toward change-related semantics without modifying any backbone weights. Finally, low-rank adaptation (LoRA) is applied to both the attention projections and MLP layers in the last four blocks of the DINOv3 backbone, with a rank of 24, $\alpha$ of 48, and dropout of 0.1. This enables task-specific tuning with only 0.97M trainable parameters.
LoRA hu2022lora is a technique that maintains the 1.1 billion-parameter DINOv3 backbone frozen while inserting trainable rank-$r$ matrices into the last four blocks, specifically the attention projections and MLPs. This method allows for task-specific tuning without modifying the original backbone weights.
For any pre-trained weight $\mathbf{W}_{0}\in\mathbb{R}^{d\times k}$, the forward pass is modified as follows:
$\mathbf{h}=\mathbf{W}_{0}\mathbf{x}+\frac{\alpha}{r}\mathbf{B}\mathbf{A}\mathbf{x},$ (2)
where $\mathbf{B}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times k}$ are the trainable matrices, and only $\mathbf{B}$ and $\mathbf{A}$ are updated.
The rank $r$ controls the expressiveness of the model. A larger $r$ provides more basis vectors and finer corrections, but the parameter count grows approximately as $2dr$. The scaling factor $\alpha$ sets the effective step-size. An $\alpha\approx 2r$ keeps the initial update variance close to full fine-tuning. If $\alpha\ll r$, the adapter is suppressed, and if $\alpha\gg r$, it risks causing unstable gradients.
3.3 Multi-Scale Cross-Attention Difference Module
To further enhance spatio-temporal information extraction from bi-temporal remote sensing images, we propose the multi-scale cross-attention difference (MSCAD) module as a core component of our framework. As illustrated in Figure 3, the MSCAD module proceeds in three conceptual steps: cross-scale alignment, temporal-difference modeling, and discrepancy aggregation.
First, multi-level features $C_{3}$, $C_{4}$, $C_{5}$ are projected to a common 256-D space via $1\times 1$ conv, up-sampled to $C_{2}$ resolution, and fused by a channel-wise attention gate by
$\mathbf{A}=\mathrm{Softmax}\bigl(\mathrm{Conv}_{1\times 1}([\tilde{\mathbf{C}}_{3};\tilde{\mathbf{C}}_{4};\tilde{\mathbf{C}}_{5}])\bigr),$ (3)
and
$\mathbf{C}_{\mathrm{fused}}=\sum_{i=2}^{4}A_{i}\odot\tilde{\mathbf{C}}_{i}.$ (4)
Next, the concatenated tensor $\mathbf{F}_{\mathrm{cat}}$ == $\mathbf{C}_{\mathrm{fused}}^{t_{1}};\mathbf{C}_{\mathrm{fused}}^{t_{2}}]$ is fed to a lightweight stack of $3\times 3$ conv, BN, GELU and $1\times 1$ conv to model non-linear temporal discrepancies and produce an adaptive difference feature $\mathbf{D}_{\mathrm{ada}}$.
Finally, the direct difference $\mathbf{D}_{\mathrm{dir}}=|\mathbf{C}^{t_{1}}-\mathbf{C}^{t_{2}}|$ and the adaptive difference $\mathbf{D}_{\mathrm{ada}}$ are concatenated and compressed by $1\times 1$ conv, BN and GELU, yielding the multi-scale change feature:
$\mathbf{D}_{\mathrm{out}}=\mathrm{GELU}\bigl(\mathrm{BN}\bigl(\mathrm{Conv}_{1\times 1}([\mathbf{D}_{\mathrm{dir}};\mathbf{D}_{\mathrm{ada}}])\bigr)\bigr).$ (5)
3.4 Context-Aware Feature Enhancement Decoder
To refine the coarse change cues produced by the MSCAD module, we design a context-aware feature enhancement decoder, which contains two lightweight yet effective sub-blocks. First, a three-layer depthwise-separable residual stack (Figure 3) enriches spatial context via
$\mathbf{Y}=\mathbf{X}+\mathrm{BN}\bigl(\mathrm{GELU}(\mathrm{DWConv}_{3\times 3}(\mathrm{Conv}_{3\times 3}(\mathbf{X})))\bigr),$ (6)
where $\quad\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, followed by a $1\times 1$ projection that recalibrates channel statistics, sharpening thin change edges and recovering detail lost during up-sampling. Next, an element-wise sigmoid gate re-weights the refined difference map as
$\mathbf{G}=\sigma(\mathbf{Y}),$ (7)
and we have:
$\mathbf{D}_{\mathrm{out}}=\mathbf{G}\odot\mathbf{Y},$ (8)
[…]
3.5 Loss Function and Optimisation
To effectively train the proposed multi-scale cross-attention difference siamese networ (MC-DiSNet), We minimise a composite loss
$\mathcal{L}=0.4\,\mathcal{L}_{\text{Focal}}+0.3\,\mathcal{L}_{\text{Dice}}+0.3\,\mathcal{L}_{\text{Lovász}},$ (9)
whose three constituents are detailed below.
Focal loss. lin2017focal counter-acts foreground/background imbalance by down-weighting easy pixels and up-weighting hard ones. With class-balanced weight $\alpha$ and focusing factor $\gamma=3$,
$\mathcal{L}_{\text{Focal}}=-\alpha(1-p_{t})^{\gamma}\log p_{t},$ (10)
where $p_{t}$ denotes the predicted probability of the ground-truth class.
Dice loss. li2020dice Dice loss maximises the overlap between prediction and ground-truth masks. Let $y\in\{0,1\}$ and $\hat{y}\in[0,1]$ be the ground-truth and predicted probabilities, respectively, then
$\mathcal{L}_{\text{Dice}}=1-\frac{2\sum y\hat{y}+\varepsilon}{\sum y^{2}+\sum\hat{y}^{2}+\varepsilon},$ (11)
with $\varepsilon=10^{-5}$ for numerical stability.
Lovász-Softmax loss. berman2018lovasz Lovász-Softmax loss is a convex surrogate of the IoU that directly optimises the intersection-over-union measure. For a mini-batch of $N$ pixels and $C$ classes,
$\mathcal{L}_{\text{Lovász}}=\sum_{c=1}^{C}\frac{1}{|\mathcal{Y}_{c}|}\sum_{i\in\mathcal{Y}_{c}}\Delta_{\text{IoU}}(p_{ic}),$ (12)
where $\mathcal{Y}_{c}$ indexes pixels of class $c$ and $\Delta_{\text{IoU}}$ is the Lovász extension of the IoU loss.
4 Experiments
4.1 Experimental Setup
4.1.1 Datasets Description
In this section, we describe all the datasets we used, the evaluation metrics, and the experimental settings. There are three datasets to evaluate the proposed change detection method, one of which is a new dataset that we provide.
We release the Gaza-Change dataset, the first open-source bi-temporal benchmark for the Gaza Strip ($31.4^{\circ}N$-$31.6^{\circ}N$, $34.4^{\circ}N$-$34.6^{\circ}N$). It contains 922 precisely co-registered image pairs $512\times 512$ acquired by the Beijing-2 satellite ($3.2m$ GSD) during 2023-2024. The imagery covers nine major urban areas, including Khan Yunis and Rafah. In line with our novel paradigm that focuses annotations exclusively on altered areas rather than full bi-temporal semantics, each pixel is meticulously annotated with one of six fine-grained change categories: building damage, new building, new camp, farmland damage, greenhouse damage, and new greenhouse. The dataset is randomly split into 554 training, 184 validation, and 184 test pairs. As shown in Figure 4, $T1$ and $T2$ show the bi-temporal images, and the label image illustrates the change labels with different colors. Beyond filling the data gap for semantic change detection task in conflict zones, Gaza-Change facilitates multi-dimensional humanitarian assessment and promotes scenario-specific algorithm development. Its six-class taxonomy characterizes the nature of each transition rather than a simple binary change, offering richer evidence for policy-making and resource allocation.
The SEmantic Change detectiON Dataset (SECOND)is a large-scale, pixel-wise annotated benchmark designed for SCD in high-resolution aerial imagery yang2021asymmetric. It consists of 4662 bi-temporal image pairs ($512\times 512$ pixels), collected over the Chinese cities of Hangzhou, Chengdu, and Shanghai using multiple platforms and sensors. Each pair was independently labeled by remote sensing experts to ensure high annotation fidelity. In contrast to our targeted annotation paradigm, SECOND provides exhaustive semantic labels for both time points, distinguishing six fundamental land-cover classes: non-vegetated ground surface, tree, low vegetation, water, buildings, and playgrounds.
Table 1 summarizes the specific information of two aforementioned datasets. It is important to emphasize that while both the Gaza-Change and SECOND datasets are annotated with semantic labels. To validate our proposed task of semantic change detection, the labels in the SECOND dataset were also processed to retain semantic information exclusively for the changed areas.
Table 1: Main information of two datasets.
Dataset Patch Size Resolution Types of Changes Train Val Test
Gaza-Change $\times512$ 3.2m 6 554 184 184
SECOND 512 $\times$512 0.5-3m 30 2968 - 1694
Figure 4: Examples of the proposed Gaza-Change, six distinct colors to encode change categories: red highlights “Building Damage", green marks “New Building", blue indicates “New Camp", yellow denotes "Farmland Damage", purple signals "Greenhouse Damage", whereas cyan represents "New Greenhouse".
4.1.2 Evaluation Metrics
In the domain of change detection, several traditional metrics are commonly used to evaluate the performance of models. These include overall accuracy (OA), precision (P), recall (R), mean intersection-over-union (mIoU), and F1 score. Each metric provides a different perspective on the model’s performance, and they are crucial for understanding the effectiveness of change detection algorithms. Mathematically, let $\mathrm{TP}_{c}$ denote the number of true positives for class $c$, and $\mathrm{FP}_{c}$ denote the number of false positives for class $c$. Then $C$ is the total number of classes, the key metrics are as follows:
$\mathrm{OA}=\frac{\sum_{i=0}^{N}q_{ii}}{\sum_{i=0}^{N}\sum_{j=0}^{N}q_{ij}}$ (13)
$\mathrm{Precision}_{c}=\frac{\mathrm{TP}_{c}}{\mathrm{TP}_{c}+\mathrm{FP}_{c}}$ (14)
$\mathrm{Recall}_{c}=\frac{\mathrm{TP}_{c}}{\mathrm{TP}_{c}+\mathrm{FN}_{c}}$ (15)
$\mathrm{mIoU}=\frac{1}{C}\sum_{c=1}^{C}\frac{\mathrm{TP}_{c}}{\mathrm{TP}_{c}+\mathrm{FP}_{c}+\mathrm{FN}_{c}}$ (16)
$\mathrm{F1}_{c}=2\cdot\frac{\mathrm{Precision}_{c}\cdot\mathrm{Recall}_{c}}{\mathrm{Precision}_{c}+\mathrm{Recall}_{c}},$ (17)
where $q_{ii}$ represents the number of correctly classified pixels for class $i$, and $q_{ij}$ represents the number of pixels predicted as class $i$ but actually belonging to class $j$.
4.1.3 Experimental Details
All experiments were conducted on a single Nvidia RTX3090Ti GPU with 24 GB VRAM. Every model was trained for 200 epochs with a batch size of 4. A common augmentation pipeline (random horizontal, vertical flip, and random rotation) was applied to all training samples. All compared methods and our proposal share the identical optimizer recipe: AdamW wen2025fantastic with learning rate $3\times 10^{-4}$, weight decay $0.01$, and $\beta=(0.9,0.999)$. The learning rate is scheduled by cosine annealing with warm restarts ($T_{0}=30$, $T_{\text{mult}}=2$, $\eta_{\min}=1\times 10^{-7}$).
To accelerate the learning of newly introduced lightweight modules, we apply parameter-wise LR scaling (ours only) by multiplying the base learning rate as follows: backbone.adapters and backbone.prompt_tokens by 20, decoderhead by 8, and backbone (frozen) by 0. All implementations are built on PyTorch 2.1.1 and trained from scratch under the above unified configuration.
4.2 Results
We report the results achieved by the model checkpoint that performs best on the test set. Table 2 summarizes the quantitative comparison between MC-DiSNet and nine representative methods, including both CNN-based (SNUNet fang2021snunet) and Transformer-based (BIT chen2021remote, ChangeFormer bandara2022transformer) baselines, as well as the most recent ChangeMamba chen2024changemamba.
  • SNUNet fang2021snunet. Employs densely-connected nested UNet architecture with channel attention modules to enhance feature reuse efficiency and improve detail preservation in change detection tasks.
  • LGPNet liu2021building. Utilizes local-global pyramid networks to capture multi-scale building features, combined with attention mechanisms to enhance boundary accuracy in urban building change detection.
  • BIT chen2021remote. A transformer-based framework for bi-temporal image processing, using semantic tokens to efficiently model change regions in remote sensing imagery
  • ChangeFormer bandara2022transformer. A pure transformer architecture for change detection that extracts multi-scale global features through hierarchical feature pyramids, significantly improving long-range dependency modeling.
  • SARASNet chen2023saras. Incorporate relation-aware, scale-aware, and interaction modules to enhance the spatial and scale perception capabilities of the Siamese network.
  • STNet ma2023stnet. A remote sensing change detection network that introduces cross-temporal gating and cross-scale attention mechanisms for spatiotemporal feature fusion
  • USSFCNet 10081023. Unified spatial-spectral frequency channel network that utilizes frequency domain transformations to mine deep feature representations, improving change detection robustness in complex scenarios.
  • DDLNeT ma2024ddlnet. A dual-domain learning network for remote sensing change detection that enhances change features in the frequency domain using discrete cosine transform while recovering spatial details.
  • ChangeMamba chen2024changemamba. Lightweight architecture based on state space models, achieving efficient long-sequence modeling through selective scanning mechanisms while balancing global receptive fields and computational efficiency.
  • Rsmamba liu2024rscama. Visual state space model designed for remote sensing images, combining convolutional locality advantages with state space sequence modeling capabilities to optimize global context capture.
  • CDxLstm wu2025cdxlstm. Hybrid architecture integrating dilated convolutions with LSTM, enhancing detection performance for gradual changes through multi-scale temporal feature extraction.
Figure 5: Example results on Gaza-change dataset.
As reported in Table 2, our MC-DiSNet achieves 86.10% precision, 60.98% recall, 55.16% mIoU, and 69.25% F1 score. The proposed MC-DiSNet exhibits a highly lightweight architecture, containing only 0.97 million parameters, which is 42 times fewer than ChangeFormer and 12 times fewer than BIT. Our framework outperforms the previous best method, DDLNet, by a large margin of 12.0% in F1 score and 12.9% in mIoU, underscoring its significant advantages in accuracy and efficiency.
As shown in Figure 5 and Figure 6, we provide representative visual comparisons of different change detection methods on sample cases from the Gaza-Change dataset. These qualitative results clearly demonstrate our method’s advantages. As shown in the first row, for scenarios with relatively distinct categories and simpler contexts, most methods can accurately identify semantic changes to a certain extent. The second row demonstrates that for relatively dense building damage, methods like CDxLstm and ChangeFormer exhibit fewer false positives but suffer from more missed detections. In comparison, our method achieves a better balance between overall false positives and false negatives. From the third and fourth rows, it can be observed that for agricultural areas and weed fields with highly similar shapes and colors, most models struggle to achieve accurate identification. Traditional convolutional methods such as USSCFNet, SNUNet, and LGPNet tend to produce discontinuous patchy detection results, whereas our method generates the fewest false alarms while maintaining detection continuity.
Figure 6: Qualitative comparison on the Gaza-Change dataset, with color coding: true positives (white), true negatives (black), false positives (red), and false negatives (green).
Table 2: Comparison with mainstream change detection methods on the Gaza-change dataset.
Method Params (M) P(%) R(%) mIoU(%) F1(%) Best Epoch
USSFCNET 1.50 46.66 48.75 34.43 47.20 51
BIT 11.90 57.45 54.90 42.13 56.15 65
SNUNet 12.00 51.44 51.42 36.50 50.35 49
DDLNET 13.80 56.61 59.96 42.23 57.22 34
STNet 14.60 59.15 54.96 42.81 56.64 53
CDxLstm 16.20 56.96 47.29 36.84 51.25 83
ChangeFormer 41.10 59.06 46.52 36.62 50.50 88
Rsmamba 52.00 58.43 42.81 36.45 49.40 66
LGPNet 71.00 55.69 55.69 48.64 54.64 64
Ours 0.97 86.10 60.98 55.16 69.25 29
To further verify the effectiveness and generalizability of the proposed method, we conduct experiments on an extra SCD datasets: SECOND. Quantitative results are summarised in Table 3. Obviously, our method achieves an mIoU of 27.74% and an F1-score of 40.59%, surpassing the strongest competitor (BIT) by 6.87 percentage points in mIoU and 8.39 percentage points in F1-score. This consistent superiority on an external dataset demonstrates that the proposed modules do not overfit the primary training data and generalize effectively to other remote sensing scenarios with varying imaging conditions and annotation granularity.
Table 3: Performance comparison on SECOND datasets.
Method SECOND
mIoU (%) F1 (%)
BIT 20.87 32.20
SNUNet 12.88 20.85
SARASNet 16.38 25.34
USSFCNet 12.03 17.80
ChangeFormer 15.48 28.30
CDxLstm 15.19 27.25
RSMamba 12.54 20.09
LGPNet 12.42 19.37
Ours 27.74 40.59
As demonstrated on our own dataset as well as on the additional benchmarks, the proposed approach consistently delivers state-of-the-art results not only on small-scale sets (hundreds of image pairs) but also on large-scale collections (tens of thousands of samples). This robust performance across dramatically different data volumes indicates strong generalization capability and practical deployment potential for diverse remote-sensing scenarios.
4.3 Ablation Study
To systematically evaluate the contributions of each proposed component, we conduct two groups of ablation experiments: (1) module-level ablation on the multi-scale attention mechanism, and (2) parameter-level ablation on the LoRA-specific hyperparameters.
4.3.1 Module-level Ablation
MSCAD fuses multi-scale C3-C5 features into a shared 256-D space and re-weights them with a channel-wise gate, allowing the model to exploit both fine detail and high-level semantics. We evaluated the multi-scale attention mechanism in MSCAD, as shown in Table 4.
Table 4: Module-level ablation results on the validation set. Best scores are in bold.
Module P(%) R(%) mIoU (%) F1 (%)
w/ attention 86.10 60.98 55.16 69.25
w/o attention 89.37 54.35 50.28 64.99
Compared with the baseline without attention, the method incorporating attention mechanism achieved a 4.88% improvement in mIoU and a 4.26% increase in F1 score. These results demonstrate that the multi-scale attention mechanism significantly enhances the model’s feature extraction capability. Although precision shows a slight decrease, the substantial improvements in both recall and F1 score confirm the effectiveness of multi-scale features in detecting small targets.
4.3.2 Parameter-level Ablation
Table 5 compares different combinations of rank $r$ and scaling factor $\alpha$. As shown in Table.5, the best trade-off between accuracy and parameter budget is achieved at r=24 and $\alpha$=48, where the model reaches 69.25% F1 and 55.16% mIoU with only 0.968 M additional trainable weights.
Table 5: Ablation on LoRA rank $r$ and scaling factor $\alpha$.
$r$ $\alpha$ Params(M) mIoU (%) F1 (%)
4 4 0.16 54.30 68.30
4 8 0.16 51.70 66.21
8 8 0.32 52.75 67.43
8 16 0.32 51.08 65.47
16 16 0.64 52.68 66.08
16 32 0.64 53.81 68.25
24 24 0.97 55.08 68.96
24 48 0.97 55.16 69.25
32 32 1.29 54.27 68.89
32 64 1.29 52.83 66.94
Increasing r from 8 to 24 yields a +3.78% F1 improvement, while further growing r to 32 drops F1 by 2.31% and almost doubles the parameters, revealing a clear saturation point.
Maintaining $\alpha/r\approx 2$ consistently outperforms $\alpha/r=1$, confirming that a slightly larger scaling factor helps the low-rank adapter capture stronger change representations. Consequently, we adopt $lora_{r}=24$ and $lora{\alpha}=48$ as the default LoRA configuration in all subsequent experiments.
5 Discussion
Our study makes three primary contributions to conflict zone monitoring through remote sensing. First, we introduce a novel dataset specifically designed for assessing land cover changes in conflict-affected areas, categorized into six critical classes: building damage, new building, new camp, farmland damage, greenhouse damage, and new greenhouse. This dataset addresses a significant gap in available benchmarks for humanitarian damage assessment.
Second, we propose a new paradigm that directly annotates changed areas rather than separately labeling bi-temporal semantic regions. This approach, which we term Change Semantic Detection (CSD), effectively extends traditional Binary Change Detection (BCD) to semantic change analysis while avoiding error accumulation from intermediate processing stages. The CSD framework substantially reduces both annotation workload and complexity, enabling rapid damage assessment in conflict scenarios.
Third, to address the challenge of limited feature representation in small targets, we develop the Multi-scale Cross-attention Difference Siamese Network (MC-DiSNet), which leverages pre-trained foundation models to enhance feature extraction capability. As reported in Table 6, our method achieves promising performance on several critical damage categories, with F1-scores of 74.64% for greenhouse damage, 64.54% for new greenhouse, and 76.85% for building damage.
Table 6: Quantitative results of our MC-DiSNet on the proposed Gaza-change dataset under the CSD task, reporting per-category performance metrics for all six change types.
Class OA(%) P(%) R(%) F1(%) IoU(%)
building damage 95.55 98.92 75.26 85.48 74.64
new building 99.52 73.67 59.32 65.72 48.94
new camp 99.57 70.21 44.10 54.17 37.15
farmland damage 97.69 99.00 28.95 44.79 28.86
greenhouse damage 99.74 89.24 69.98 78.45 64.54
new greenhouse 99.86 85.58 88.28 86.91 76.85
The moderate performance on new building and new camp (45.21% and 52.67% respectively) likely stems from their relatively small spatial extent in the imagery. Notably, farmland damage shows the lowest performance (28.86%), potentially due to the significant feature distribution differences between agricultural areas and other structural damage categories, suggesting that our model may require further adaptation to effectively capture diverse land cover characteristics.
These findings highlight both the promise and challenges of automated damage assessment in conflict zones, while demonstrating the potential of our CSD paradigm and MC-DiSNet architecture to advance the field of humanitarian remote sensing.
6 Conclusion
In this work, we have introduced a comprehensive framework for semantic change detection in conflict zones through three key contributions. First, we presented a novel dataset specifically designed for assessing conflict-induced land changes, featuring six critical damage and construction categories. Second, we proposed the CSD paradigm, which directly focuses on changed areas to enable efficient damage assessment while significantly reducing annotation costs. Third, we developed the MC-DiSNet that effectively leverages pre-trained vision foundation models to enhance feature representation for small targets. Our experimental results demonstrate that the proposed method achieves strong performance on both Gaza-change and SECOND datasets.
This work takes a significant step toward automated conflict damage assessment using remote sensing imagery. We believe our contributions in dataset creation, task formulation, and methodology development will facilitate future research in humanitarian remote sensing and emergency response.
Data availability
The Gaza-change dataset presented in this study is available to qualified researchers upon reasonable request.
Acknowledgments
The authors extend sincere appreciation to the DeepSeek large language model for its valuable assistance in text polishing and language refinement during the writing of this paper.

### v2

@@
 It is worth emphasizing that our annotations only include semantic pixels of changed areas.
-Unlike conventional semantic change detection (SCD), our approach eliminates the need for large-scale semantic annotations of bi-temporal images, instead focusing directly on the changed regions.
-We term this new task change semantic detection (CSD).
-The CSD task represents a direct extension of binary change detection (BCD).
-Due to the limited spatial extent of semantic regions, it presents greater challenges than traditional SCD tasks.
-We evaluated our method under the CSD framework on both the Gaza-Change and SECOND datasets.
-Experimental results demonstrate that our proposed approach effectively addresses the CSD task, and its outstanding performance paves the way for practical applications in rapid damage assessment across conflict zones.
+Unlike conventional semantic change detection (SCD), our approach eliminates the need for large-scale semantic annotations of bi-temporal images, instead focusing directly on the changed regions, which terms multi-class change detection (MCD).
+We evaluated our method on the Gaza-Change and two classical datasets: SECOND and Landsat-SCD datasets.
+Experimental results demonstrate that our proposed approach effectively addresses the MCD task, and its outstanding performance paves the way for practical applications in rapid damage assessment across conflict zones.
 keywords:
@@
 Nevertheless, a fundamental challenge persists: these data-hungry deep models require vast amounts of meticulously annotated data, which is notoriously difficult and expensive to obtain for remote sensing applications.
-To address the aforementioned challenges, we propose a new paradigm termed change semantic detection (CSD).
-Distinct from conventional semantic change detection (SCD), our approach eliminates the need for annotating entire semantic regions, instead focusing solely on change masks.
+To address the aforementioned challenges, we introduce the multi-class change detection (MSD).
+Distinct from conventional semantic change detection (SCD), MCD eliminates the need for annotating entire semantic regions, instead focusing solely on change masks.
 This framework represents a direct extension of binary change detection (BCD).
@@
 This necessitates substantial improvements in the model’s capability to extract features from small target areas.
-Figure 1: Evolution of change detection paradigms: (a) Binary Change Detection (BCD), (b) Semantic Change Detection (SCD), and (c) our proposed Change Semantic Detection (CSD).
-Based on the above analysis, we identify four major challenges in framing conflict-induced damage assessment as a CSD task: (1) Inherent data scarcity: Limited by the geographical extent of conflict zones and the number of destroyed areas available for training.
-(2) Small target regions: CSD focuses exclusively on damaged areas, resulting in minimal semantic region coverage.
+Figure 1: Evolution of change detection paradigms: (a) Binary Change Detection (BCD), (b) Semantic Change Detection (SCD), and (c) Multi-Class Change Detection (MCD).
+Based on the above analysis, we identify four major challenges in framing conflict-induced damage assessment as a MCD task: (1) Inherent data scarcity: Limited by the geographical extent of conflict zones and the number of destroyed areas available for training.
+(2) Small target regions: MCD focuses exclusively on damaged areas, resulting in minimal semantic region coverage.
 (3) Subtle and ambiguous changes: Infrastructure damage in conflict zones varies significantly in severity and extent, particularly making minor damage difficult to detect.
@@
 We argue that leveraging their rich, pre-trained representations is key to overcoming data scarcity and recognizing subtle semantic changes.
-In this paper, we propose a novel DINOv3-driven siamese network for CSD.
+In this paper, we propose a novel DINOv3-driven siamese network for MCD.
 Specifically, we adopt the DINOv3 simeoni2025dinov3 model pre-trained on satellite data, with ConvNeXt liu2022convnet as its main backbone architecture, which helps reduce the distribution discrepancy between the pre-training data and the actual application data.
@@
 To our knowledge, this is the first change detection study specifically focused on conflict area assessment.
-  • We propose the novel change semantic detection (CSD) paradigm that fundamentally shifts from exhaustive bi-temporal semantic annotation to focused labeling of changed semantic regions.
+  • We introduce the multi-class change detection (MCD) paradigm for damage assessment that fundamentally shifts from exhaustive bi-temporal semantic annotation to focused labeling of changed semantic regions.
 This strategic simplification significantly reduces annotation complexity and human labor.
-  • Extensive experiments show that our method achieves state-of-the-art performance not only on our proposed Gaza dataset but also on the public benchmarks, SECOND yang2021asymmetric, demonstrating its superior robustness and generalization capability.
+  • Extensive experiments show that our method achieves state-of-the-art performance not only on our proposed Gaza dataset but also on the public benchmarks, SECOND yang2021asymmetric and Landsat-SCD yuan2022transformer, demonstrating its superior robustness and generalization capability.
 Figure 2: The panoramic remote sensing satellite image of the Gaza Strip.
@@
 3.2 Enhanced DINOv3 Backbone
-To maximize the feature extraction capability of the DINOv3 backbone while balancing computational cost and accuracy for the proposed change semantic detection (CSD) task, we introduce an enhanced DINOv3 encoder.
+To maximize the feature extraction capability of the DINOv3 backbone while balancing computational cost and accuracy for the multi-class change detection (MCD) task, we introduce an enhanced DINOv3 encoder.
 This encoder keeps the original ConvNeXt-Tiny weights frozen and enriches multi-scale features C3-C5 through three lightweight plug-ins.
@@
 $\mathbf{C}_{\mathrm{fused}}=\sum_{i=2}^{4}A_{i}\odot\tilde{\mathbf{C}}_{i}.$ (4)
-Next, the concatenated tensor $\mathbf{F}_{\mathrm{cat}}$ == $\mathbf{C}_{\mathrm{fused}}^{t_{1}};\mathbf{C}_{\mathrm{fused}}^{t_{2}}]$ is fed to a lightweight stack of $3\times 3$ conv, BN, GELU and $1\times 1$ conv to model non-linear temporal discrepancies and produce an adaptive difference feature $\mathbf{D}_{\mathrm{ada}}$.
-Finally, the direct difference $\mathbf{D}_{\mathrm{dir}}=|\mathbf{C}^{t_{1}}-\mathbf{C}^{t_{2}}|$ and the adaptive difference $\mathbf{D}_{\mathrm{ada}}$ are concatenated and compressed by $1\times 1$ conv, BN and GELU, yielding the multi-scale change feature:
+Next, the concatenated tensor $\mathbf{F}_{\mathrm{cat}}$ == $\mathbf{C}_{\mathrm{fused}}^{t_{1}};\mathbf{C}_{\mathrm{fused}}^{t_{2}}]$ is fed to a lightweight stack of $3\times 3$ conv, BN, GELU and $1\times 1$ conv to model non-linear temporal discrepancies and produce an adaptive difference feature $\mathbf{D}_{\mathrm{ada}}$ (diff_ada in Figure 3).
+Finally, the direct difference $\mathbf{D}_{\mathrm{dir}}=|\mathbf{C}^{t_{1}}-\mathbf{C}^{t_{2}}|$ (diff_abs in Figure 3) and the adaptive difference $\mathbf{D}_{\mathrm{ada}}$ are concatenated and compressed by $1\times 1$ conv, BN and GELU, yielding the multi-scale change feature diff_aggergator:
 $\mathbf{D}_{\mathrm{out}}=\mathrm{GELU}\bigl(\mathrm{BN}\bigl(\mathrm{conv}_{1\times 1}([\mathbf{D}_{\mathrm{dir}};\mathbf{D}_{\mathrm{ada}}])\bigr)\bigr).$ (5)
-3.4 Context-Aware Feature Enhancement Decoder
-To refine the coarse change cues produced by the MSCAD module, we design a context-aware feature enhancement decoder, which contains two lightweight yet effective sub-blocks.
+3.4 Decoder
+To refine the coarse change cues produced by the MSCAD module, we design a enhancement decoder, which contains two lightweight yet effective sub-blocks.
 First, a three-layer depthwise-separable residual stack (Figure 3) enriches spatial context via
@@
 and we have:
-$\mathbf{D}_{\mathrm{out}}=\mathbf{G}\odot\mathbf{Y},$ (8)
+$\mathbf{F}_{\mathrm{out}}=\mathbf{G}\odot\mathbf{Y},$ (8)
 […]
@@
 In contrast to our targeted annotation paradigm, SECOND provides exhaustive semantic labels for both time points, distinguishing six fundamental land-cover classes: non-vegetated ground surface, tree, low vegetation, water, buildings, and playgrounds.
-Table 1 summarizes the specific information of two aforementioned datasets.
-It is important to emphasize that while both the Gaza-Change and SECOND datasets are annotated with semantic labels.
-To validate our proposed task of semantic change detection, the labels in the SECOND dataset were also processed to retain semantic information exclusively for the changed areas.
-Table 1: Main information of two datasets.
+Landsat-SCD yuan2022transformer provides ten change types and is constructed based on Landsat-like images acquired between 1990 and 2020 in Tumushuke, Xinjiang.
+The region is located along the Belt and Road Economic Belt, adjacent to the Taklimakan Desert, and features a fragile ecological environment.
+The data source consists of Landsat series imagery with a spatial resolution of 30 meters, which offers relatively high spatiotemporal and spectral resolution as well as good data accessibility.
+Table 1 summarizes the specific information of the three aforementioned datasets.
+It is important to emphasize that while both the Gaza-Change, SECOND, and Landsat-SCD datasets are annotated with semantic labels.
+To validate the MCD task, the labels in the SECOND and Landsat-SCD dataset were also processed to retain semantic information exclusively for the changed areas.
+Table 1: Main information of three datasets.
 Dataset Patch Size Resolution Types of Changes Train Val Test
@@
 SECOND 512 $\times$512 0.5-3m 30 2968 - 1694
+Landsat-SCD 416 $\times$ 416 30m 9 1431 477 477
 Figure 4: Examples of the proposed Gaza-Change, six distinct colors to encode change categories: red highlights “Building Damage", green marks “New Building", blue indicates “New Camp", yellow denotes "Farmland Damage", purple signals "Greenhouse Damage", whereas cyan represents "New Greenhouse".
@@
 All implementations are built on PyTorch 2.1.1 and trained from scratch under the above unified configuration.
-4.2 Results
-We report the results achieved by the model checkpoint that performs best on the test set.
-Table 2 summarizes the quantitative comparison between MC-DiSNet and nine representative methods, including both CNN-based (SNUNet fang2021snunet) and Transformer-based (BIT chen2021remote, ChangeFormer bandara2022transformer) baselines, as well as the most recent ChangeMamba chen2024changemamba.
+All compared methods are described as follows:
   • SNUNet fang2021snunet.
@@
   • CDxLstm wu2025cdxlstm. Hybrid architecture integrating dilated convolutions with LSTM, enhancing detection performance for gradual changes through multi-scale temporal feature extraction.
+4.2 Main Results
+Result of Caza-change dataset.
+We report the results achieved by the model checkpoint that performs best on the test set.
+Table 2 summarizes the quantitative comparison between MC-DiSNet and nine representative methods.
 Figure 5: Example results on Gaza-change dataset.
@@
 Table 2: Comparison with mainstream change detection methods on the Gaza-change dataset.
-Method Params (M) P(%) R(%) mIoU(%) F1(%) Best Epoch
-USSFCNET 1.50 46.66 48.75 34.43 47.20 51
-BIT 11.90 57.45 54.90 42.13 56.15 65
-SNUNet 12.00 51.44 51.42 36.50 50.35 49
-DDLNET 13.80 56.61 59.96 42.23 57.22 34
-STNet 14.60 59.15 54.96 42.81 56.64 53
-CDxLstm 16.20 56.96 47.29 36.84 51.25 83
-ChangeFormer 41.10 59.06 46.52 36.62 50.50 88
-Rsmamba 52.00 58.43 42.81 36.45 49.40 66
-LGPNet 71.00 55.69 55.69 48.64 54.64 64
-Ours 0.97 86.10 60.98 55.16 69.25 29
-To further verify the effectiveness and generalizability of the proposed method, we conduct experiments on an extra SCD datasets: SECOND.
+Method Trainable Params (M) P(%) R(%) mIoU(%) F1(%) Best Epoch
+USSFCNET 1.5 46.66 48.75 34.43 47.20 51
+BIT 11.9 57.45 54.90 42.13 56.15 65
+SNUNet 12.0 51.44 51.42 36.50 50.35 49
+DDLNET 13.8 56.61 59.96 42.23 57.22 34
+STNet 14.6 59.15 54.96 42.81 56.64 53
+CDxLstm 16.2 56.96 47.29 36.84 51.25 83
+ChangeFormer 41.1 59.06 46.52 36.62 50.50 88
+Rsmamba 52.0 58.43 42.81 36.45 49.40 66
+LGPNet 71.0 55.69 55.69 48.64 54.64 64
+Ours 18.2 86.10 60.98 55.16 69.25 29
+Results of SECOND and Landsat-SCD dataset.
+To further verify the effectiveness and generalizability of the proposed method, we conduct experiments on the two extra SCD datasets: SECOND and Landsat-SCD.
 Quantitative results are summarised in Table 3.
 Obviously, our method achieves an mIoU of 27.74% and an F1-score of 40.59%, surpassing the strongest competitor (BIT) by 6.87 percentage points in mIoU and 8.39 percentage points in F1-score.
-This consistent superiority on an external dataset demonstrates that the proposed modules do not overfit the primary training data and generalize effectively to other remote sensing scenarios with varying imaging conditions and annotation granularity.
-Table 3: Performance comparison on SECOND datasets.
-Method SECOND
-mIoU (%) F1 (%)
-BIT 20.87 32.20
-SNUNet 12.88 20.85
-SARASNet 16.38 25.34
-USSFCNet 12.03 17.80
-ChangeFormer 15.48 28.30
-CDxLstm 15.19 27.25
-RSMamba 12.54 20.09
-LGPNet 12.42 19.37
-Ours 27.74 40.59
-As demonstrated on our own dataset as well as on the additional benchmarks, the proposed approach consistently delivers state-of-the-art results not only on small-scale sets (hundreds of image pairs) but also on large-scale collections (tens of thousands of samples).
+On the large-scale Landsat-SCD dataset, our method achieves a mIoU of 68.79% and an F1-score of 80.69%, significantly outperforming the previous state-of-the-art LGPNet (61.48% mIoU and 75.45% F1).
+This represents notable absolute gains of +7.31% in mIoU and +5.24% in F1, corresponding to relative improvements of 11.9% and 6.9%, respectively.
+This consistent superiority on two external datasets demonstrates that the proposed modules do not overfit the primary training data and generalize effectively to other remote sensing scenarios with varying imaging conditions and annotation granularity.
+Table 3: Performance comparison on SECOND and Landsat-SCD datasets.
+Method SECOND Landsat-SCD
+mIoU (%) F1 (%) mIoU (%) F1 (%)
+BIT 20.87 32.20 60.13 74.33
+SNUNet 12.88 20.85 42.07 56.28
+USSFCNet 12.03 17.80 41.05 55.14
+ChangeFormer 15.48 28.30 57.09 71.45
+CDxLstm 15.19 27.25 47.93 62.89
+RSMamba 12.54 20.09 55.41 70.00
+LGPNet 12.42 19.37 61.48 75.45
+Ours 27.74 40.59 68.79 80.69
+Figure 7: Example results on Landsat-SCD dataset.
+Figure 8: Qualitative comparison on the Landsat-SCD dataset, with color coding: true positives (white), true negatives (black), false positives (red), and false negatives (green).
+We also present visual detection results on the Landsat-SCD dataset in Figure 7 and Figure 8.
+As shown in the third row of Figure 7, MC-DiSNet produces a continuous and uniformly wide delineation of the narrow, curved change belt, whereas results from other methods exhibit fragmentation or over-expansion.
+The first row of Figure 8 demonstrates that our result aligns closely with the ground truth (GT) without spurious speckles, while other methods show extensive green false negatives (FN).
+However, it is important to note that while our method significantly reduces under-detection, it introduces a certain amount of false positives (red pixels), leading to an overall reddish hue in some areas.
+This observation underscores a broader challenge: effective change detection models often require targeted adaptation to the specific characteristics of different remote sensing datasets.
+Overall, As demonstrated on our own dataset as well as on the additional benchmarks, the proposed approach consistently delivers state-of-the-art results not only on small-scale sets (hundreds of image pairs) but also on large-scale collections (tens of thousands of samples).
 This robust performance across dramatically different data volumes indicates strong generalization capability and practical deployment potential for diverse remote-sensing scenarios.
-4.3 Ablation Study
-To systematically evaluate the contributions of each proposed component, we conduct two groups of ablation experiments: (1) module-level ablation on the multi-scale attention mechanism, and (2) parameter-level ablation on the LoRA-specific hyperparameters.
-4.3.1 Module-level Ablation
-MSCAD fuses multi-scale C3-C5 features into a shared 256-D space and re-weights them with a channel-wise gate, allowing the model to exploit both fine detail and high-level semantics.
-We evaluated the multi-scale attention mechanism in MSCAD, as shown in Table 4.
 Table 4: Module-level ablation results on the validation set. Best scores are in bold.
-Module P(%) R(%) mIoU (%) F1 (%)
-w/ attention 86.10 60.98 55.16 69.25
-w/o attention 89.37 54.35 50.28 64.99
-Compared with the baseline without attention, the method incorporating attention mechanism achieved a 4.88% improvement in mIoU and a 4.26% increase in F1 score.
-These results demonstrate that the multi-scale attention mechanism significantly enhances the model’s feature extraction capability.
-Although precision shows a slight decrease, the substantial improvements in both recall and F1 score confirm the effectiveness of multi-scale features in detecting small targets.
-4.3.2 Parameter-level Ablation
-Table 5 compares different combinations of rank $r$ and scaling factor $\alpha$.
-As shown in Table.5, the best trade-off between accuracy and parameter budget is achieved at r=24 and $\alpha$=48, where the model reaches 69.25% F1 and 55.16% mIoU with only 0.968 M additional trainable weights.
+Module MS-att diff-ada diff-agg att P(%) R(%) mIoU (%) F1 (%)
+Encoder 89.37 54.35 50.28 64.99
+79.26 49.20 44.25 58.36
+78.77 50.50 46.10 59.09
+Decoder 83.39 60.24 53.81 67.26
+Base 86.10 60.98 55.16 69.25
+Table 5: Comparison with different backbones in encoder of our proposed MC-DiSNet on the Gaza-change dataset.
+Method P(%) R(%) mIoU(%) F1(%) Params (M) Trainable Params (M)
+Dinov3-tiny (ours) 86.10 60.98 55.16 69.25 29.09 1.78
+Dinov3-smal1 77.52 65.35 62.03 70.71 53.06 3.25
+ResNet-18 51.56 40.13 32.45 42.99 14.95 0.68
+ResNet-50 59.37 39.07 34.03 44.48 25.76 1.55
+SAM-base 90.61 59.95 56.46 71.65 95.39 1.36
+Swinformer-tiny 73.71 63.47 58.05 67.92 29.58 1.62
 Table 6: Ablation on LoRA rank $r$ and scaling factor $\alpha$.
@@
 32 64 1.29 52.83 66.94
+Table 7: Quantitative results of our MC-DiSNet on the proposed Gaza-change dataset under the MCD task, reporting per-category performance metrics for all six change types.
+Class OA(%) P(%) R(%) F1(%) IoU(%)
+building damage 95.55 98.92 75.26 85.48 74.64
+new building 99.52 73.67 59.32 65.72 48.94
+new camp 99.57 70.21 44.10 54.17 37.15
+farmland damage 97.69 99.00 28.95 44.79 28.86
+greenhouse damage 99.74 89.24 69.98 78.45 64.54
+new greenhouse 99.86 85.58 88.28 86.91 76.85
+4.3 Ablation Study
+To systematically evaluate the contributions of each proposed component, we conduct two groups of ablation experiments: (1) module-level ablation on the multi-scale attention mechanism, and (2) parameter-level ablation on the LoRA-specific hyperparameters.
+4.3.1 Module-level Ablation
+We conduct ablation studies on key components of MC-DiSNet: (a) the multi-scale attention (MS-att) in the encoder, (b) the two feature fusion modules, diff_Ada and diff_agg, and (c) the attention mechanism (Att) in the decoder.
+Table 4 reports the comparison results.
+To systematically validate the necessity and synergistic effects of the core components in MC-DiSNet, we conducted module ablation experiments on the validation set, with results presented in Table 4.
+The full model with all components (Base) achieved the best performance (55.16% mIoU, 69.25% F1).
+Removing components individually led to performance degradation to varying degrees: Removing the multi-scale attention (MS-att) caused mIoU to drop by 4.88% and F1 by 4.26%, indicating its crucial role in multi-scale discriminative feature selection.
+Removing the decoder attention (att) reduced mIoU to 53.81% and lowered the recall rate, confirming its core function in refining the final change map.
+Removing diff_agg resulted in one of the most severe performance drops (mIoU 46.10%, F1 59.09%), highlighting the irreplaceability of cross-level fusion of difference features.
+Finally, completely removing diff_ada led to the largest degradation (mIoU 44.25%, F1 58.36%), demonstrating its key role in suppressing false-change noise.
+[…]
+4.3.2 Parameter-level Ablation
+Table 6 compares different combinations of rank $r$ and scaling factor $\alpha$.
+As shown in Table.6, the best trade-off between accuracy and parameter budget is achieved at r=24 and $\alpha$=48, where the model reaches 69.25% F1 and 55.16% mIoU with only 0.97 M additional trainable weights.
 Increasing r from 8 to 24 yields a +3.78% F1 improvement, while further growing r to 32 drops F1 by 2.31% and almost doubles the parameters, revealing a clear saturation point.
@@
 This dataset addresses a significant gap in available benchmarks for humanitarian damage assessment.
-Second, we propose a new paradigm that directly annotates changed areas rather than separately labeling bi-temporal semantic regions.
-This approach, which we term Change Semantic Detection (CSD), effectively extends traditional Binary Change Detection (BCD) to semantic change analysis while avoiding error accumulation from intermediate processing stages.
-The CSD framework substantially reduces both annotation workload and complexity, enabling rapid damage assessment in conflict scenarios.
+Second, we introduce the MCD task that directly annotates changed areas rather than separately labeling bi-temporal semantic regions.
+This approach effectively extends traditional Binary Change Detection (BCD) to semantic change analysis while avoiding error accumulation from intermediate processing stages.
+The MCD framework substantially reduces both annotation workload and complexity, enabling rapid damage assessment in conflict scenarios.
 Third, to address the challenge of limited feature representation in small targets, we develop the Multi-scale Cross-attention Difference Siamese Network (MC-DiSNet), which leverages pre-trained foundation models to enhance feature extraction capability.
 As reported in Table 7, our method achieves promising performance on several critical damage categories, with F1-scores of 74.64% for greenhouse damage, 64.54% for new greenhouse, and 76.85% for building damage.
-Table 6: Quantitative results of our MC-DiSNet on the proposed Gaza-change dataset under the CSD task, reporting per-category performance metrics for all six change types.
-Class OA(%) P(%) R(%) F1(%) IoU(%)
-building damage 95.55 98.92 75.26 85.48 74.64
-new building 99.52 73.67 59.32 65.72 48.94
-new camp 99.57 70.21 44.10 54.17 37.15
-farmland damage 97.69 99.00 28.95 44.79 28.86
-greenhouse damage 99.74 89.24 69.98 78.45 64.54
-new greenhouse 99.86 85.58 88.28 86.91 76.85
 The moderate performance on new building and new camp (45.21% and 52.67% respectively) likely stems from their relatively small spatial extent in the imagery.
 Notably, farmland damage shows the lowest performance (28.86%), potentially due to the significant feature distribution differences between agricultural areas and other structural damage categories, suggesting that our model may require further adaptation to effectively capture diverse land cover characteristics.
-These findings highlight both the promise and challenges of automated damage assessment in conflict zones, while demonstrating the potential of our CSD paradigm and MC-DiSNet architecture to advance the field of humanitarian remote sensing.
+These findings highlight both the promise and challenges of automated damage assessment in conflict zones, while demonstrating the potential of the MCD paradigm and MC-DiSNet architecture to advance the field of humanitarian remote sensing.
 6 Conclusion
@@
 First, we presented a novel dataset specifically designed for assessing conflict-induced land changes, featuring six critical damage and construction categories.
-Second, we proposed the CSD paradigm, which directly focuses on changed areas to enable efficient damage assessment while significantly reducing annotation costs.
+Second, we introduced the MCD paradigm, which directly focuses on changed areas to enable efficient damage assessment while significantly reducing annotation costs.
 Third, we developed the MC-DiSNet that effectively leverages pre-trained vision foundation models to enhance feature representation for small targets.
-Our experimental results demonstrate that the proposed method achieves strong performance on both Gaza-change and SECOND datasets.
+Our experimental results demonstrate that the proposed method achieves strong performance on the Gaza-change, SECOND, and Landsat-SCD datasets.
 This work takes a significant step toward automated conflict damage assessment using remote sensing imagery.